A factory worker places 86 newly created circuits on a shelf to be checked for quality. of these 6 will not work correctly. Suppose that she is asked to randomly select 2 circuits, without replacement. what is the chance that both circuits she selects will be defective? approximate to the nearest ten thousandth.
Answer by stanbon(26259)

A factory worker places 86 newly created circuits on a shelf to be checked for quality. of these 6 will not work correctly. Suppose that she is asked to randomly select 2 circuits, without replacement. what is the chance that both circuits she selects will be defective? approximate to the nearest ten thousandth.
---------------------------
# of ways to pick 2 defective: 6C2
# of ways to pick 2 or the 86: 86C2
---------------
Ans: 6C2/86C2 = 15/3655 = 0.0041
===============================
Cheers,
Stan H.